Randy is four times as old as Marie. In 12yr, Marie's age will be half of Randy's age. How old are they now?
Answer by ptaylor(2198)

Let x=Marie's age; x+12=Marie's age in 12 years
Then 4x=Randy's age; 4x+12=Randy's age in 12 years
Now we are told that:
x+12=1/2(4x+12) simplify
x+12=2x+6 subtract x and also 6 from each side
x-x+12-6=2x-x+6-6
6=x or
x=6 years old-------------Marie's age now
4x=4*6=24 years old--------Randy's age now
CK
6+12=1/2(24+12)
18=1/2(36)
18=18
Hope this helps--ptaylor